You are an economic consultant for the aqua-agriculture industry. You are in a position to provide some economic advice to a business in this sector. You have estimated the supply equation using regression estimation techniques. The variables that you have included are Pa (the price of the product itself); Pb and Pc (the prices of related goods in production; Pw reflects wages in the industry; Pi is the price of other inputs (raw materials). Your research indicates that the equation is as follows: Qs = -10 + 5Pa + 4Pb - 3Pc - 10Pw - 5Pi (a) Explain the difference between Goods B and C. How are these goods related differently to Good A? (b) What is the relationship between wages and the price of good A? Explain your answer. (c) If the price of Good B is $20, and the price of Good C is $15, while the wage rate is $15, and the price of other inputs is $10, what is the P intercept of the supply equation? Show the exact calculations for your answer. (d) Suppose that the wage rate decreases to $10. What is the impact on the supply curve? Explain your answer and show the exact calculations for your answer.
Answer #1

A)we need to make indirect supply curve to know difference between good. B and c.

Pa=0.2Q-0.8pb+0.6pc+2pw+2

Look at the sign associated with good b and c price.

There is positive sign with good c price ,which tells if good c price get lower , seller will charge lower price for same QUANTITY,means there is a competition between good c and good a ,so the good c is a substitute of good a.

On other hand negitive sign with good b price which tells good b price gets lower ,seller will charge higher price for same QUANTITY,means there is collusion between good b and a ,so the good c and a are complementary goods.

B) there is positive sign with wage means Increase in wage induce seller to increase price for same QUANTITY and supply decrease.

C)Pa=0.2Q-0.8pb+0.6pc+2pw+pi+5

At p intercept, Q=0

pa=0-0.8*20+0.6*15+2*15+10+5=-16+9+30+10+5=38

D)as stated in b). There is a negitive relationship between wage and supply

At wage =20

Q=5pb-295

At wage =0

Q=-10+5pb-4*20+3*15-10*10-5*10=-10+5pb-80+45-100-50

Q=5pb-195